OpenStudy (anonymous):

plse answer this question................. three vertices of a square are P(-1,-9) , Q(3,-1) and R(-5,3) . plot these points on graph . and find the coordinate of missing vertex......

OpenStudy (rajee_sam):

why are YOU asking this question???

OpenStudy (anonymous):

i forgot how to find the missing (fourth ) vertex

OpenStudy (radar):



OpenStudy (radar):

It is a square (in spite of what it looks like lol) and all sides are equal in lenght and the corners are right angles (the lines are perpendicular)

OpenStudy (anonymous):

hmm, i got the same figure

OpenStudy (radar):

I would first find the slope of the line RQ. delta y over delta x, giving us 4/-8 = -1/2 we know that line SP has to be parallel so the slope for that line is also -1/2. Then the equation for that line SP is (y + 9)/(x + 1) = -1/2 y= -(1/2)x - 19/2 Now calculate the equation for the line RS slope = 8/4 = 2, line equation is: y = 2x + 13 You now have a system of equations, 2 equations, 2 unknowns. y= -(1/2)x - 19/2 y= 2x + 13 y=y -(1/2x)-19/2 =2x +13 multiply thru by 2 -x-19=4x + 26 -5x = 45 x=-9 y=2(-9) +13 y=-5 The 4th vertice is located at (-9, -5)

OpenStudy (anonymous):

@radar thanks bro

OpenStudy (radar):

You are welcome, there are I'm sure many other ways to have found that, like using the distance formula etc., but I went that way.

OpenStudy (radar):

Good luck in your studies, I will sign off for now.